OpenStudy (anonymous):

please help! this question is making me fraustrated! medals rewarded:) what is the least common multiple, LCM, of the two polynomials? (x+3) and (x^2-9)

OpenStudy (anonymous):

x² - 9 = x² - 3² = ( x - 3 ) ( x+ 3) Can you continue?

OpenStudy (anonymous):

nope that perfect thankyou!

OpenStudy (anonymous):

can you do this one for me quick?

OpenStudy (anonymous):

4/m+2+2/m-2

OpenStudy (anonymous):

I can't recognize which one is denominator ?

OpenStudy (anonymous):

m+2 and m-2 is the denominator:)

OpenStudy (anonymous):

(4)/(m+2)+(2)/(m-2)

OpenStudy (anonymous):

4/ ( m+2) + 2/ ( m -2) =[ 4 ( m-2) + 2 ( m + 2) ] / (m² - 4 )

OpenStudy (anonymous):

Do you understantd the common denominator process?

OpenStudy (anonymous):

ya kind the choices are a..6/(m+2)(m-2) b.6m-4/(m+2)(m-2) c.6m+12/(m+2)(m-2)

OpenStudy (anonymous):

which one would it be?

OpenStudy (anonymous):

Just simply the denominator, try it!

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

Try again!

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

You can't distribute the denominator out?

OpenStudy (anonymous):

it b isnt it lol sorry I know what I did wrong lol